Problem 7-50 (Static) Assigning Costs: Missing Data (LO 7-2, 3)

The following T-accounts represent September activity.

Required:

Compute the missing amounts indicated by the letters (a) through (i).


Materials Inventory
BB (9/1) 8,000
(a) 4,300
(b)
EB (9/30) 9,700
Work-in-Process Inventory
BB (9/1) 22,300
180,500 (e)
121,000
94,000
EB (9/30) 17,700
Finished Goods Inventory
BB (9/1) 14,200
(e) (f)
EB (9/30) (g)
Cost of Goods Sold
402,800
Applied Overhead Control
(d)
Manufacturing Overhead Control
121,000
4,300
36,200
31,600
3,200
Wages Payable
124,300
162,000 (c)
36,200
119,500 EB (9/30)
Accumulated Depreciation—Plant & Equipment
204,100 BB (9/1)
(h)
235,700 EB (9/30)
Accounts Payable—Material Suppliers
100,000
Prepaid Expenses
BB(9/1) 24,300
(i)
EB(9/30) 21,100
Answer #1

b. $180500
a. $180500+4300+8000-1700 = $191100
c. $121000
d. $94000
e. $22300+180500+121000+94000-17700 = $400100
f. $402800
g. $14200+400100-402800 = $11500
h. $31600
i. $3200